OpenStudy (anonymous):

Which table represents the increasing linear function with the greatest unit rate? x y 2 16 5 10 x y 2 14 6 12 x y 2 -24 5 -15 x y 2 -12 6 -16 x y 2 -19 6 -17

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@Compassionate

OpenStudy (michele_laino):

here you have to choose one ordered pair from each table. Foe example from first table I choose \((2,16)\). In this case the unit rate is: \(16/2=8\). Now please do the same with the remaining two tables of values

OpenStudy (michele_laino):

oops.. For* example

OpenStudy (michele_laino):

please wait a moment

OpenStudy (anonymous):

so 5/10 = 2

OpenStudy (michele_laino):

sorry, we have to apply another procedure. We have to write the equation for each of your three group of data

OpenStudy (anonymous):

so the answer is B

OpenStudy (michele_laino):

for example, I start with the data collected in first table, namely: (2,14) and (6,12) I write the henral equation of a linear function like below: \(y=ax+b\) then I replace the coordinates above, so I get this algebraic system: \[\left\{ \begin{gathered} 14 = 2a + b \hfill \\ 12 = 6a + b \hfill \\ \end{gathered} \right.\] Now, I have to find the values of \(a\) and \(b\)

OpenStudy (michele_laino):

oops.. the general* formula

OpenStudy (michele_laino):

If I solve such algebraic system, I find: \(a=-1/2\) and \(b=15\), so the linear function which models the data collected into your first table is: \(y=-x/2 +15\), and the unit rate is \(-1/2\) Now, please do the same, using data of the other two tables